OpenStudy (anonymous):

Solve: -2 < 5 x + 4 le 1

OpenStudy (anonymous):

What do you mean by "4 le 1"?

OpenStudy (anonymous):

less than or equal to

OpenStudy (anonymous):

\[-2 < 5x+4 \le 1\] -4 -4 -4 \[-6 < 5x \le -3\] Divide everything by 5 \[\frac{ -6 }{ 5 } < x \le \frac{ -3 }{ 3 }\]

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

You're welcome!